OpenStudy (anonymous):

6x^2=-4x in order to solve the equation x would equal 0 and 2/3??

OpenStudy (anonymous):

6x^2=-4x 6x^2+4x = 0 3x^2+2x =0 solve with quadratic formual

OpenStudy (anonymous):

i have no clue what quadratic formula is

OpenStudy (anonymous):

you can factorize if you don't wanna use quadratic. you will get x(6x - 4) = 0 => x = 0 or 6x - 4 = 0 6x = 4 x = 4/6 simplified : 2/3